# LLM Multi-Agent Systems: Challenges and Open Problems

### v2

@@ 1 Introduction
Multi-agent systems enhance the capabilities of single LLM agents by leveraging collaborations among agents and their specialized abilities (Talebirad & Nadiri, 2023; Zhang et al., 2023a; Park et al., 2023; Li et al., 2023; Jinxin et al., 2023). It utilizing collaboration and coordination among agents to execute tasks that are beyond the capability of any individual agent. In multi-agent systems, each agent is equipped with distinctive capabilities and roles, collaborating towards the fulfillment of some common objectives. Such collaboration, characterized by activities such as debate and reflection, has proven particularly effective for tasks requiring deep thought and innovation. Recent works include simulating interactive environments (Park et al., 2023; Jinxin et al., 2023), role-playing (Li et al., 2023), reasoning (Du et al., 2023; Liang et al., 2023), demonstrating the huge potential of multi-agent systems in handling complex real-world scenarios.
While existing works have demonstrated the impressive capabilities of multi-agent systems, the potential for advanced multi-agent systems far exceeds the progress made to date. A large number of existing works focus on devising planning strategies within a single agent by breaking down the tasks into smaller, more manageable tasks (Chen et al., 2022; Ziqi & Lu, 2023; Yao et al., 2023; Long, 2023; Besta et al., 2023; Wang et al., 2022b). Yet, multi-agent systems involve agents of various specializations and more complex interactions and layered context information, which poses challenges to the designing of the work flow as well as the whole system. Also, existing literature pays limited attention to memory storage, while memory plays a critical role in collaborations between agents. It enables agents to access to some common sense, aligning context with their tasks, and further, learn from past work flows and adapt their strategies accordingly.
@@ 2.1 Structure of Multi-agent Systems
Hierarchical Structure. Hierarchical structures (Gronauer & Diepold, 2022; Ahilan & Dayan, 2019) typically consists of a leader and one or multiple followers; see Figure 1(b). The leader’s role is to guide or plan, while the followers respond or execute based on the leader’s instructions. Hierarchical structures are prevalent in scenarios where coordinated efforts directed by a central authority are essential. Multi-agent systems that explore Stackelberg games (Von Stackelberg, 2010; Conitzer & Sandholm, 2006) fall into this category (Harris et al., 2023). This type of game is distinguished by this leadership-followership dynamic and the sequential nature of decision-making. Agents make decisions in a sequential order, where the leader player first generate an output (e.g., instructions) then the follower players take an action based on the leader’s instruction.
@@
Dynamic Structure. Dynamic structures mean that the states of the multi-agent system, e.g., the role of agents, their relations, and the number of agents in the multi-agent system, may change (Talebirad & Nadiri, 2023) over time. As an example,  (Talebirad & Nadiri, 2023) enables addition and removal of agents to make the system to suit the tasks at hand. A multi-agent system may also be contextually adaptive, with the interaction patterns inside the system being modified based on internal system states or external factors, such as contexts. Agents in such systems can dynamically reconfigure their roles and relationships in response to changing conditions.
@@ 3.1 Global Planning
Game Theory. Game theory provides a well-structured framework for understanding strategic interactions in multi-agent systems, particularly for systems that involve complex interactions among agents such as debates or discussions. A crucial concept in game theory is equilibrium, e.g., Nash Equilibrium (Kreps, 1989) and Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer & Sandholm, 2006), that describes a state where, given the strategies of others, no agent benefits from unilaterally changing their strategy. Game theory has been applied in multi-agent systems, especially Stackelberg equilibrium (Gerstgrasser & Parkes, 2023; Harris et al., 2023), as the structure of Stackelberg equilibrium contains is a leader agent and multiple follower agents, and such hierarchical architectures are wildely considered in multi-agent systems. (Gerstgrasser & Parkes, 2023) designs a general multi-agent framework to identify Stackelberg Equilibrium in Markov games, and (Harris et al., 2023) extend the Stackelberg model to allow agents to consider external context information, such as traffic and weather, etc. However, some problems are still challenging in multi-agent systems, such as defining an appropriate payoff structure for both the collective strategy and individual agents based on the context of the overall tasks, and efficiently achieving equilibrium states. These unresolved issues highlight the ongoing need for refinement in the application of game theory to complex multi-agent scenarios.
@@ 3.2 Single-Agent Task Decomposition
Task decomposition in a single agent involves generating a series of intermediate reasoning steps to complete a task or arrive at an answer. This process can be represented as transforming direct input-output (inputoutputdelimited-$\langle\text{input}\rightarrow\text{output}\rangle$⟨ input → output ⟩) mappings into the inputrationaloutputdelimited-$\langle\text{input}\rightarrow\text{rational}\rightarrow\text{output}\rangle$⟨ input → rational → output ⟩ mappings (Wei et al., 2022; Zhang et al., 2023b). Task composition can be of different formats, as follows.
@@
iv) Table-of-Thoughts (Tab-CoT) (Ziqi & Lu, 2023) that utilize a tabular-format for reasoning, enabling the complex reasoning process to be explicitly modelled in a highly structured manner.
@@
Consistency in Objectives. In multi-agent systems, consistency in objectives is maintained across various levels, i.e., from overall goals down to individual agent tasks and their decomposed tasks. Each agent must understand and effectively utilize the layered contexts while ensuring its task and the decomposed sub-tasks to remain aligned with the overall goals. (Harris et al., 2023) extends the Stackelberg model (Von Stackelberg, 2010; Conitzer & Sandholm, 2006) to enable agents to incorporate external context information, such as context (or insights) provided by other agents. However, aligning the complex context with the decomposed tasks during reasoning remains unresolved.
@@ 4 Agent Memory and Information Retrieval
The memory in single-LLM agent systems refers to the agent’s ability to record, manage, and utilize data, such as past historical queries and some external data sources, to help inference and enhance decision-making and reasoning (Yao et al., 2023; Park et al., 2023; Li & Qiu, 2023; Wang et al., 2023; Guo et al., 2023). While the memory in a single-LLM agent system primarily focuses on internal data management and utilization, a multi-agent system requires agents to work collaboratively to complete some tasks, necessitating the individual memory capabilities of each agent as well as a sophisticated mechanism for sharing, integrating, and managing information across agents, thus poses challenges to memory and information retrieval.
@@ 5.2 Blockchain Nodes as Agents
Smart Contract Management and Optimization. Smart contracts are programs that execute the terms of a contract between a buyer and a seller in a blockchain system. The codes are fixed, and are self-executed when predetermined conditions are met. Multi-agent systems can automate and optimize the execution of smart contracts with more flexible terms and even dynamic external information from users. Agents can negotiate contract terms on behalf of their users, manage contract execution, and even optimize gas fees (in the context of Ethereum (Wood et al., 2014). The agents can analyze context information , such as past actions and pre-defined criteria, and utilize the information with flexibility. Such negotiations can also utilize game theory, such as Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer & Sandholm, 2006) when there is a leader negotiator and Nash Equilibrium (Kreps, 1989) when no leader exists.

### v3

@@ 1 Introduction
Multi-agent systems enhance the capabilities of single LLM agents by leveraging collaborations among agents and their specialized abilities (Talebirad and Nadiri, 2023; Zhang et al., 2023a; Park et al., 2023; Li et al., 2023; Jinxin et al., 2023). It utilizing collaboration and coordination among agents to execute tasks that are beyond the capability of any individual agent. In multi-agent systems, each agent is equipped with distinctive capabilities and roles, collaborating towards the fulfillment of some common objectives. Such collaboration, characterized by activities such as debate and reflection, has proven particularly effective for tasks requiring deep thought and innovation. Recent works include simulating interactive environments (Park et al., 2023; Jinxin et al., 2023), role-playing (Li et al., 2023), reasoning (Du et al., 2023; Liang et al., 2023), demonstrating the huge potential of multi-agent systems in handling complex real-world scenarios.
While existing works have demonstrated the impressive capabilities of multi-agent systems, the potential for advanced multi-agent systems far exceeds the progress made to date. A large number of existing works focus on devising planning strategies within a single agent by breaking down the tasks into smaller, more manageable tasks (Chen et al., 2022; Ziqi and Lu, 2023; Yao et al., 2023; Long, 2023; Besta et al., 2023; Wang et al., 2022b). Yet, multi-agent systems involve agents of various specializations and more complex interactions and layered context information, which poses challenges to the designing of the work flow as well as the whole system. Also, existing literature pays limited attention to memory storage, while memory plays a critical role in collaborations between agents. It enables agents to access to some common sense, aligning context with their tasks, and further, learn from past work flows and adapt their strategies accordingly.
@@ 2.1 Structure of Multi-agent Systems
Hierarchical Structure. Hierarchical structures (Gronauer and Diepold, 2022; Ahilan and Dayan, 2019) typically consists of a leader and one or multiple followers; see Figure 1(b). The leader’s role is to guide or plan, while the followers respond or execute based on the leader’s instructions. Hierarchical structures are prevalent in scenarios where coordinated efforts directed by a central authority are essential. Multi-agent systems that explore Stackelberg games (Von Stackelberg, 2010; Conitzer and Sandholm, 2006) fall into this category (Harris et al., 2023). This type of game is distinguished by this leadership-followership dynamic and the sequential nature of decision-making. Agents make decisions in a sequential order, where the leader player first generate an output (e.g., instructions) then the follower players take an action based on the leader’s instruction.
@@
Dynamic Structure. Dynamic structures mean that the states of the multi-agent system, e.g., the role of agents, their relations, and the number of agents in the multi-agent system, may change (Talebirad and Nadiri, 2023) over time. As an example,  (Talebirad and Nadiri, 2023) enables addition and removal of agents to make the system to suit the tasks at hand. A multi-agent system may also be contextually adaptive, with the interaction patterns inside the system being modified based on internal system states or external factors, such as contexts. Agents in such systems can dynamically reconfigure their roles and relationships in response to changing conditions.
@@ 3.1 Global Planning
Game Theory. Game theory provides a well-structured framework for understanding strategic interactions in multi-agent systems, particularly for systems that involve complex interactions among agents such as debates or discussions. A crucial concept in game theory is equilibrium, e.g., Nash Equilibrium (Kreps, 1989) and Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer and Sandholm, 2006), that describes a state where, given the strategies of others, no agent benefits from unilaterally changing their strategy. Game theory has been applied in multi-agent systems, especially Stackelberg equilibrium (Gerstgrasser and Parkes, 2023; Harris et al., 2023), as the structure of Stackelberg equilibrium contains is a leader agent and multiple follower agents, and such hierarchical architectures are wildely considered in multi-agent systems. (Gerstgrasser and Parkes, 2023) designs a general multi-agent framework to identify Stackelberg Equilibrium in Markov games, and (Harris et al., 2023) extend the Stackelberg model to allow agents to consider external context information, such as traffic and weather, etc. However, some problems are still challenging in multi-agent systems, such as defining an appropriate payoff structure for both the collective strategy and individual agents based on the context of the overall tasks, and efficiently achieving equilibrium states. These unresolved issues highlight the ongoing need for refinement in the application of game theory to complex multi-agent scenarios.
@@ 3.2 Single-Agent Task Decomposition
Task decomposition in a single agent involves generating a series of intermediate reasoning steps to complete a task or arrive at an answer. This process can be represented as transforming direct input-output ($\langle\text{input}\rightarrow\text{output}\rangle$) mappings into the $\langle\text{input}\rightarrow\text{rational}\rightarrow\text{output}\rangle$ mappings (Wei et al., 2022; Zhang et al., 2023b). Task composition can be of different formats, as follows.
@@
iv) Table-of-Thoughts (Tab-CoT) (Ziqi and Lu, 2023) that utilize a tabular-format for reasoning, enabling the complex reasoning process to be explicitly modelled in a highly structured manner.
@@
Consistency in Objectives. In multi-agent systems, consistency in objectives is maintained across various levels, i.e., from overall goals down to individual agent tasks and their decomposed tasks. Each agent must understand and effectively utilize the layered contexts while ensuring its task and the decomposed sub-tasks to remain aligned with the overall goals. (Harris et al., 2023) extends the Stackelberg model (Von Stackelberg, 2010; Conitzer and Sandholm, 2006) to enable agents to incorporate external context information, such as context (or insights) provided by other agents. However, aligning the complex context with the decomposed tasks during reasoning remains unresolved.
@@ 4 Agent Memory and Information Retrieval
The memory in single-LLM agent systems refers to the agent’s ability to record, manage, and utilize data, such as past historical queries and some external data sources, to help inference and enhance decision-making and reasoning (Yao et al., 2023; Park et al., 2023; Li and Qiu, 2023; Wang et al., 2023; Guo et al., 2023). While the memory in a single-LLM agent system primarily focuses on internal data management and utilization, a multi-agent system requires agents to work collaboratively to complete some tasks, necessitating the individual memory capabilities of each agent as well as a sophisticated mechanism for sharing, integrating, and managing information across agents, thus poses challenges to memory and information retrieval.
@@ 5.2 Blockchain Nodes as Agents
Smart Contract Management and Optimization. Smart contracts are programs that execute the terms of a contract between a buyer and a seller in a blockchain system. The codes are fixed, and are self-executed when predetermined conditions are met. Multi-agent systems can automate and optimize the execution of smart contracts with more flexible terms and even dynamic external information from users. Agents can negotiate contract terms on behalf of their users, manage contract execution, and even optimize gas fees (in the context of Ethereum (Wood and others, 2014). The agents can analyze context information , such as past actions and pre-defined criteria, and utilize the information with flexibility. Such negotiations can also utilize game theory, such as Stackelberg Equilibrium (Von Stackelberg, 2010; Conitzer and Sandholm, 2006) when there is a leader negotiator and Nash Equilibrium (Kreps, 1989) when no leader exists.